OpenStudy (anonymous):

In a certain county, the number of Charter Schools is eight less than twice the number of Alternative Schools. We know that there are 36 Charter Schools in the county. How many Alternative Schools are in the county?

OpenStudy (aripotta):

2x - 8 = 36. solve for x and you'll get the answer

OpenStudy (anonymous):

HOW DO YOU DO THAT ?

OpenStudy (aripotta):

ok, so x represents the amount of alternative schools and n represents the amount of charter schools. if they said that the number of charter schools is 8 less than twice the number of alternative schools, then 2x - 8 = n. they also said that there's 36 charter schools, so 2x - 8 = 36

OpenStudy (aripotta):

to solve for x, you have to follow SADMEP, the opposite of the order of operations

OpenStudy (aripotta):

do you know what to do first?

OpenStudy (anonymous):

NO

OpenStudy (aripotta):

ok, you have add 8 to both sides first

OpenStudy (anonymous):

No

OpenStudy (aripotta):

yes